2. The director of a secretarial school believes that its graduates can type more than75

words per minute. A random sample of 12 graduates has been found to have an

average of 77.2 words per minute with a standard deviation of 7.9 words per minute

in a typing test. Using the 0.05 level of significance, test the claim of the director.


1
Expert's answer
2022-06-10T14:14:25-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le75"

"H_1:\mu>75"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=11" and the critical value for aright-tailed test is "t_c =1.795885."

The rejection region for thisright-tailed test is "R = \{t:t>1.795885\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{77.2-75}{7.9/\sqrt{12}}=0.9647"


Since it is observed that "t=0.9647<1.795885=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed, "df=11" degrees of freedom, "t=0.9647" is "p=0.177712," and since "p=0.177712>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 75, at the "\alpha = 0.05" significance level.